OpenStudy (lovelyanna):

Find the slope and y-intercept of the line. y = x – 5 A. slope: –5, y-intercept: 1 B. slope: –1, y-intercept: 5 C. slope: 5, y-intercept: –1 D. slope: 1, y-intercept: –5 I think either C or D. Please help and explain. Thanks!

OpenStudy (lovelyanna):

I already know that, can you just tell me which is right and explain please?

OpenStudy (precal):

y=x-5 y=mx+b m is your slope value and b is your y intercept

OpenStudy (lovelyanna):

Ya, I know that.. I'm still confused on how to solve it.

OpenStudy (precal):



OpenStudy (precal):

what number is in front of x? that is your slope value what number does not have an x? that is your y intercept You really are not solving anything, you are just identifying.

OpenStudy (lovelyanna):

So.. D?

OpenStudy (precal):

yes

OpenStudy (lovelyanna):

Oh, ok. Thank you!

OpenStudy (precal):

yw

OpenStudy (lovelyanna):

Do you want a medal? I'm new and don't really know how to do that.